Shadi has 4.60 in quarters, dimes, and nickels. She has three more dimes than nickels and three more quarters than dimes. How many of each coin does she have?: Shadi has 4.60 in quarters, dimes, and nickels. She has three more dimes than nickels and three more quarters than dimes. How many of each coin does she have?
Answer by orca(336)

Let x be the number of dimes, then the number of nickels is x - 3, the number of quarters is x + 3.
Their total value is 10x + 5(x - 3) + 25(x + 3)
As the total value is $4.60(or 460 cents), we can set up an equation:
10x + 5(x - 3) + 25(x + 3) = 460
Solving for x, we have
10x + 5x - 15 + 25x + 75 = 460
40x = 400
x = 10
So the number of dimes is 10.
The number of nickels is x - 3 = 10 - 3 = 7.
The number of quarters is x + 3 = 10 + 3 = 13.